please help me solve this problem.
manuel traveled 4 hours non stop to mexico a total of 166 mles. he took a train part of the way which averaged 50 mph, and then took a bus the remaining distance, which avereged 30 mph. how long was manuel on the train?

Found 2 solutions by mananth, lwsshak3:
Answer by mananth(16946)

First part distance = x--------------speed =50 mph
II part will be 166-x----------------------30 mph
t=4
time first part + time II part = 4
t=d/r
x/50 + (166-x)/30 = 4
LCD = 150
multiply equation by 150
3x+5(166-x)=4*150
3x+830-5x=600
-2x=-230
/-2
x= 115 miles
time on train = 115 /50=2.3 hours

Answer by lwsshak3(11628)

manuel traveled 4 hours non stop to mexico a total of 166 mles. he took a train part of the way which averaged 50 mph, and then took a bus the remaining distance, which avereged 30 mph. how long was manuel on the train?
**
let x=miles Manuel was on the train
166-x=miles Manuel was on the bus
travel time=distance/speed
x/50+(166-x)/30=4
LCD=50*30
30x+50(166-x)=4*50*30
30x+8300-50x=6000
20x=2300
x=115 miles
Travel time=distance/speed
115/50=2.3 hrs
ans:
Manuel was on the train for 2.3 hrs